# Simultaneous Evolutionary Fits for Jupiter and Saturn Incorporating Fuzzy Cores

### v1

@@ 2 APPLE Code
where $X_{i}=\{X,Y,Z\}$ denotes the hydrogen, helium, and heavy element mass fractions, respectively, and all other symbols have standard definitions (see Sur et al., 2024). A compact core of mass $M_{c}$, using the AQUA EOS (Haldemann et al., 2020), is embedded in the center. The code supports various H-He EOS options from the EOS module of Tejada Arevalo et al. (2024a), such as SCvH95 (Saumon et al., 1995), CMS19+HG23 (Chabrier et al., 2019; Howard et al., 2023), and CD21 (Chabrier & Debras, 2021), with heavy elements mixed via the volume addition law.
@@ 4.1 Jupiter Model
Building on these results, we present an updated Jupiter model that retains the features of Tejada Arevalo et al. (2024b), while also now reproducing the gravitational moments $J_{2}$ and $J_{4}$ (Durante et al., 2020). This updated model benefits from enhanced numerical smoothing of thermodynamic derivatives in our equation of state module (Tejada Arevalo et al., 2024a). Gravitational moments are calculated using the Theory of Figures to fourth order (ToF4, Nettelmann, 2017), which also provides the moment of inertia and rotation-corrected equatorial radius. These quantities are important in the centrifugal term of the equation of hydrostatic equilibrium (eq. 1), under the assumption of solid-body rotation, to provide a second-order correction to account for model oblateness.
@@
Figure 3 illustrates the evolution of our best-fit Jupiter model. It is very similar to that found by Tejada Arevalo et al. (2024b), with slight differences due to the slightly smoother thermal derivatives employed and the desire here simultaneously to fit the lower gravitational moments. This model predicts a total heavy element mass of 42.5 M, of which 3 M resides in the core with $Z=1$. To match Jupiter’s observed $Y_{\rm atm}$, the required temperature shift for the LHR0911 miscibility curve is +410 K (compared to +300 K in Tejada Arevalo et al., 2024b) with helium rain initiating at 4 Gyr. The resulting best-fit model numbers with the corresponding percent differences (in parentheses) from the measured values are: current effective temperature $T_{\rm eff}=124.6$ K ($\sim 0.7\%$), $Y_{\rm atm}=0.236$ ($\sim 0.8\%$), $Z_{\rm atm}=3.6$ Z, and equatorial radius $=$ $R_{\rm eq}=72019.5$ km ($\sim 0.7\%$). The gravitational moments $J_{2}$ and $J_{4}$ are $14731.6\times 10^{-6}$ ($\sim 0.2\%$) and $-591.46\times 10^{-6}$ ($\sim 0.3\%$), respectively.
@@ 4.2 Saturn Model
Figure 4 illustrates the evolution of our best-fit Saturn model, which predicts a total heavy-element mass of 25 M, with 4 M located in the core. The model closely reproduces the observed values, yielding an effective temperature of 96.54 K, a deviation of 0.13% from current measurements (Li et al., 2010), an atmospheric helium fraction of $Y_{\rm atm}=0.205$, and an atmospheric metallicity of $Z_{\rm atm}\sim 4.57$ Z after 4.56 Gyr of evolution. At the current epoch, the equatorial radius is $\sim$59,551.8 km, deviating by roughly 1.18%, while the gravitational moments are $J_{2}=16365.7\times 10^{-6}$ and $J_{4}=-850.11\times 10^{-6}$, differing by 0.4% and a small margin, respectively. This model successfully matches Saturn’s observed properties, while offering insights into the sensitivity of its interior structure to variations in key parameters.
@@ 5 Discussion and Conclusions
We emphasize that we have for this work used the Ledoux condition for convective stability and have ignored the possibility of semi-convection (Chabrier & Baraffe, 2007; Rosenblum et al., 2011; Mirouh et al., 2012; Leconte & Chabrier, 2012; Wood et al., 2013; Medrano et al., 2014; Tulekeyev et al., 2024). We feel that the physics of doubly-diffusive mixing and transport is still in flux and we will leave to future work our assessment of this important topic. It may well be that the Schwarzschild condition would be better, given convective overshoot (Korre et al., 2019; Anders et al., 2022) and new insights into the devolution in some parameter regimes on short timescales of semi-convection into convection (Garaud et al., 2017; Tulekeyev et al., 2024). Moreover, the erosion of the core over time (Wilson & Militzer, 2012a, b; Moll et al., 2017; Öberg & Wordsworth, 2019), and not its production at formation, may be responsible for the current presence of a fuzzy core. Finally, there is the intriguing possibility that rapid rotation in Saturn may itself inhibit the penetration of a convective zone into its interior, thereby potentially explaining (or facilitating) the survival of a fuzzy core (Fuentes et al., 2023). This possibility, along with the potential importance of a large miscibility latent heat (Markham & Guillot, 2024), emphasizes that there is yet much to do to resolve the many important outstanding issues in Jovian planet structure and evolution. Nevertheless, we suggest that our new models, generated in the context of the new understanding of Jupiter and Saturn that has emerged in the last few years, are important first steps towards the ultimate revelation of their internal characters.

### v2

@@ 2 APPLE Code
where $X_{i}=\{X,Y,Z\}$ denotes the hydrogen, helium, and heavy element mass fractions, respectively, and all other symbols have standard definitions (see Sur et al., 2024). Heavy elements in the envelope were modeled using the AQUA EOS (Haldemann et al., 2020), while a compact core of mass $M_{c}$, is made up of 50% iron and 50% post-perovskite (MgSiO3) (Keane, 1954; Stacey & Davis, 2004; Zhang & Rogers, 2022), is embedded in the center. The code supports various H-He EOS options from the EOS module of Tejada Arevalo et al. (2024a), such as SCvH95 (Saumon et al., 1995), CMS19+HG23 (Chabrier et al., 2019; Howard et al., 2023), and CD21 (Chabrier & Debras, 2021), with heavy elements mixed via the volume addition law. The angular frequency in Equation 1 evolves over time, while conserving angular momentum. The time-dependent moment of inertia is derived from the structure using the Theory of Figures to Fourth Order (TOF4; Nettelmann, 2017). The conservation calculation assumes present-day inferred rotation rates and corresponding moment of inertia estimates (e.g., Militzer et al., 2022).
@@ 4.1 Jupiter Model
Building on these results, we present an updated Jupiter model that retains the features of Tejada Arevalo et al. (2024b), while also now approximately reproducing the gravitational moments $J_{2}$ and $J_{4}$ (Durante et al., 2020). This updated model benefits from enhanced numerical smoothing of thermodynamic derivatives in our equation of state module (Tejada Arevalo et al., 2024a). Gravitational moments are calculated using the Theory of Figures to fourth order (ToF4, Nettelmann, 2017), which also provides the moment of inertia and rotation-corrected equatorial radius. These quantities are important in the centrifugal term of the equation of hydrostatic equilibrium (eq. 1), under the assumption of solid-body rotation, to provide a second-order correction to account for model oblateness. We do not normalize either to the present-day (4.56 Gyrs.) equatorial radius, but let the code arrive naturally at values of radius, $J_{2}$, and $J_{4}$.
@@
Figure 3 illustrates the evolution of our best-fit Jupiter model. It is very similar to that found by Tejada Arevalo et al. (2024b), with slight differences due to the slightly smoother thermal derivatives employed and the desire here simultaneously to fit the lower gravitational moments. This model predicts a total heavy element mass of 42.5 M, of which 3 M resides in the core with $Z=1$. To match Jupiter’s observed $Y_{\rm atm}$, the required temperature shift for the LHR0911 miscibility curve is +410 K (compared to +300 K in Tejada Arevalo et al., 2024b) with helium rain initiating at 4 Gyr. The resulting best-fit model numbers with the corresponding percent differences (in parentheses) from the measured values are: current effective temperature $T_{\rm eff}=124.6$ K ($\sim 0.7\%$), $Y_{\rm atm}=0.236$ ($\sim 0.8\%$), $Z_{\rm atm}=3.6$ Z, and equatorial radius $=$ $R_{\rm eq}=72019.5$ km ($\sim 0.7\%$). The gravitational moments $J_{2}$ and $J_{4}$ are $14731.6\times 10^{-6}$ ($\sim 0.2\%$) and $-591.46\times 10^{-6}$ ($\mathbf{\sim 0.8\%}$), respectively. Table 1 presents a comparison of our results with the observed measurements for Jupiter.
@@ 4.2 Saturn Model
Table 1: Comparison of measured quantities and our model results for Jupiter and Saturn at 4.56 Gyr. References for the measured values: [1] Li et al. (2010), [2] Li et al. (2012), [3] Seidelmann et al. (2007), [4] von Zahn et al. (1998), [5] Achterberg & Flasar (2020); Koskinen & Guerlet (2018), [6] Guillot et al. (2023), [7] Iess et al. (2018), [8] Iess et al. (2019)
Quantity Jupiter Saturn
Measured Our Model (4.56 Gyr) Measured Our Model (4.56 Gyr)
Teff [K] 125.57 $\pm$ 0.07 [1] 124.6 96.67 [2] 96.54
Equatorial Radius [km] 71492 ±$\pm 4$ [3] 72019.5 60268 ±$\pm 4$ [3] 59,551.8
Atmospheric Helium Fraction (Yatm) 0.234 ±$\pm 0.005$ [4] 0.236 0.075-0.22 [5] 0.205
Atmospheric Metallicity (Z/Z) 1.5–5 [6] 3.6 5.0–10.0 [6] 4.6
$J_{2}$$\times 10^{-6}$ 14696.572 [7] 14731.6 16290.573 [8] 16365.7
$J_{4}$$\times 10^{-6}$586.609 [7] -591.46 -935.314 [8] -850.11
@@
Figure 4 illustrates the evolution of our best-fit Saturn model, which predicts a total heavy-element mass of 25 M, with 4 M located in the core. The model closely reproduces the observed values, yielding an effective temperature of 96.54 K, a deviation of 0.13% from current measurements (Li et al., 2010), an atmospheric helium fraction of $Y_{\rm atm}=0.205$, and an atmospheric metallicity of $Z_{\rm atm}\sim 4.57$ Z after 4.56 Gyr of evolution. At the current epoch, the equatorial radius is $\sim$59,551.8 km, deviating by roughly 1.18%, while the gravitational moments are $J_{2}=16365.7\times 10^{-6}$ and $J_{4}=-850.11\times 10^{-6}$, differing by 0.4% and a small margin, respectively. This model successfully matches Saturn’s observed properties, while offering insights into the sensitivity of its interior structure to variations in key parameters. Table 1 provides a summary of our results alongside the observed measurements for Saturn.
@@ 5 Discussion and Conclusions
We emphasize that we have for this work used the Ledoux condition for convective stability and have ignored the possibility of semi-convection (Chabrier & Baraffe, 2007; Rosenblum et al., 2011; Mirouh et al., 2012; Leconte & Chabrier, 2012; Wood et al., 2013; Medrano et al., 2014; Tulekeyev et al., 2024). We feel that the physics of doubly-diffusive mixing and transport is still in flux and we will leave to future work our assessment of this important topic. It may well be that the Schwarzschild condition would be better, given convective overshoot (Korre et al., 2019; Anders et al., 2022) and new insights into the devolution in some parameter regimes on short timescales of semi-convection into convection (Garaud et al., 2017; Tulekeyev et al., 2024). Moreover, the erosion of the core over time (Wilson & Militzer, 2012a, b; Moll et al., 2017; Öberg & Wordsworth, 2019), and not its production at formation, may be responsible for the current presence of a fuzzy core. Finally, there is the intriguing possibility that rapid rotation in Saturn may itself inhibit the penetration of a convective zone into its interior, thereby potentially explaining (or facilitating) the survival of a fuzzy core (Fuentes et al., 2023, 2024). This possibility, along with the potential importance of a large miscibility latent heat (Markham & Guillot, 2024), emphasizes that there is yet much to do to resolve the many important outstanding issues in Jovian planet structure and evolution. Nevertheless, we suggest that our new models, generated in the context of the new understanding of Jupiter and Saturn that has emerged in the last few years, are important first steps towards the ultimate revelation of their internal characters.